OpenStudy (anonymous):

Can some one help me answering these four questions?? I think I have the correct answer for the last two but im not really sure 1. how do you find the distance between two points on the coordinate plane? 2. how do you find the slope between 2 points on the coordinate plane? 3. what do you have to see as evidence to be convinced that the 2 lines are parallel? 4. what do you have to see as evidence to be convinced that 2 lines are perpendicular?

Mehek (mehek14):

what are your answers?

OpenStudy (anonymous):

Never mind I dont have the last two...I ended up loosing my answers..

Mehek (mehek14):

Parallel lines You know they're parallel when they never intersect 

Mehek (mehek14):

Perpendicular the lines intersect at a 90 degree angle 

Mehek (mehek14):

does that help for the last 2?

OpenStudy (anonymous):

Yeah but the first two im really struggling on

Mehek (mehek14):

finding the distance on a coordinate plane: distance formula

OpenStudy (anonymous):

or you could say the slope for parallel lines are exactly the same and for perpendicular the slopes are exact opposites

Mehek (mehek14):

@Chumpian perpendicular lines have to intersect at 90 degrees

OpenStudy (anonymous):

yes i know that lol

Mehek (mehek14):

2) \(\dfrac{rise}{run}\)

OpenStudy (anonymous):

Ok thank you that helps me alot

OpenStudy (anonymous):

Could you explain the distance formula really quick to me though?

Mehek (mehek14):

sure \(\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\)

Mehek (mehek14):

if you have these ordered pairs (10,0) and (5,4) \(y_2=4\\y_1=0\\x_2=5\\x_1=10\)

Mehek (mehek14):

\(\sqrt{(4-0)^2+(5-10)^2}\\ \sqrt{4^2+(-5)^2}\\ \sqrt{16 + 25} \\ \sqrt{ 41} = 6.4\)

Mehek (mehek14):

^ that's how you do it

OpenStudy (anonymous):

Thank you again

Mehek (mehek14):

yw